Light is reflected from a glass coffee table. When the angle of incidence is 60.2 degrees, the reflected light is completely polarized parallel to the surface of the glass. What is the index of refraction of the glass?
1
Expert's answer
2020-12-23T07:35:40-0500

Let's assume that the light comes to the coffee glass from the air. Then the the polarization angle (Brewster's angle) is


"\theta = \arctan (n_2)"

where "n_2" is the index of refraction of the glass. Since we know, that "\theta = 60.2\degree", the index of refraction of the glass is:


"n_2 = \tan\theta\\
n_2 = \tan 60.2\degree \approx 1.75"

Answer. 1.75.